# Entanglement measures of Majorana bound states

### v1

@@ III Results for a maximally entangled initial state
In this part, we follow the discussion of previous section and obtain the quantum correlations and state dynamic probabilities for the two systems described in Section II, taking into account a maximally entangled initial state. We remind that the first introduced setup corresponds to a Majorana system (MS), where QDs are coupled via nonlocal fermion of Majorana origin. The second setup consists of QDs coupled via a normal fermion, coined normal fermion system (NFS). Since the manipulation of QDs was shown to be a feasible task [54, 55], see also Refs. [56, 57, 58], we consider that the two QDs are in a maximally entangled initial state, while the nonlocal fermion due to MBSs is in the zero state. We note that it is necessary to consider the zero state of the nonlocal fermion in order to maintain the even parity of the system, which is the sector we investigate, see Section II. The separability of the non-local fermion with entangled QDs is justified by assuming that the topological superconductor and QD are not coupled before the evolution of the entire system starts.
@@
where |$\ket{E_{j}}$ represents the eigenstates of the MS or NFS (|$(\ket{E_{j}}\rightarrow\ket{E_{j}^{\prime}})$ given by Eqs. (4) or Eqs. (7), depending on the system under investigation. Before going further, we point out that the contribution from the eigenstates in Eq. (17) also depends on the form of the initial states. In this regard, we have that the primary contributions in the MS come from eigenstates |$\ket{E_{1}}$ and |$\ket{E_{2}}$ with energies $-\Delta$ and $\Delta$, respectively. The eigenstates |$\ket{E_{3}}$ and |$\ket{E_{4}}$ are orthogonal to the initial state, $\left\langle E_{j}|\psi(0)\right\rangle=0$, which implies that they do not contribute to the dynamics given by Eq. (17). Contrary to the MS, the dynamics of the NFS is given by the contribution of all states.
@@ III.1 Majorana system
Here, $\Delta$, $\lambda$, and $\omega$ is the lowest positive eigenvalue, the coupling between nonlocal fermion and right QD, and the energy splitting between MBSs, see Subsection 4. The coefficients $\eta$ and $\chi$ in the above state function shape the dynamics of the entanglement generations between the MBSs and the second QD. They can also be manipulated to see the type of maximally entangled state (of the Bell’s state form) that may be achieved in the evolution. Therefore, below we now compute the quantities discussed in Section II.3.
@@ III.1.1 Return probability and entanglement dynamics
By using Eqs. (15), we obtain the entanglement dynamics $E_{d}$ and the return probability $R_{p}$. For $E_{d}$ we need to specify the desired state onto which will be projected the time-evolved function |$\ket{\psi(t)}$. For this purpose, we consider |$\ket{\phi}$ describing a maximally entangled nonlocal fermion and the second QD given by
@@ III.2 Normal fermion system
In this part we carry out the same calculations as in the previous subsection but when the QDs are coupled via a normal fermion. The time-evolved state function for the initial state from Eq. (16) is here given by
@@
$\begin{split}R_{p}&=\frac{|\bar{c}_{1}(t)+\bar{c}_{4}(t)|^{2}}{2}\,,\\ E_{d}&=\frac{|\bar{c}_{1}(t)+\bar{c}_{2}(t)|^{2}}{2}\,,\\ C&=2\max(||\bar{c}_{1}(t)||\bar{c}_{4}(t)|-|\bar{c}_{2}(t)||\bar{c}_{3}(t)||)\,,\end{split}$ (26)
@@
III.3 Discussion of the state dynamics and quantum correlations
@@
Having presented the general expressions obtained of concurrence (C), discord (D), entanglement dynamics ($E_{d}$) and return probability ($R_{p}$) for the MS and the NFS with maximally entangled QDs, here we discuss their time evolution. To vizualize the time evolution, in Fig. 2 we plot C, D, $E_{d}$, and $R_{p}$ as a function of time $t$ for the two systems under consideration (MS and NFS). Fig. 2(a-d) and Fig. 2(e-h) correspond to QDs with symmetric ($\lambda=1$) and asymmetric ($\lambda\neq 1$) couplings in the MS or NFS. As discussed before, in the MS setup QDs are coupled via MBSs, while QDs are coupled via a normal fermion in the NFS. To truly exploit the inherent Majorana nonlocality, in Fig. 2 we contrast regimes at zero and finite energies.
@@
We first analyze the case of symmetric couplings and at zero energy, which corresponds to $\omega=0$ and $\lambda=1$ in Fig. 2(a,b). In the case of the MS, this regime has zero-energy MBSs and the first feature to notice is that the quantities $C$, $D$, and $R_{p}$ remain at $1$, while $E_{d}=0.25$, as shown in Fig. 2(a), see also subsection III.1. Since the initial state involves maximally entangled QDs, having a constant time evolution reveals that such a maximal entanglement is not affected. The constant time evolution can be understood by noting that the eigenstates of the Hamiltonian given by Eq. (4), except for the third eigenstate, are orthonormal to the initial state given by Eq. (17), giving a vanishing overlap that does not contribute to the dynamics as seen in Fig. 2(a). The third eigenstate of the MS is the same as the initial state [Eq. (16)] and has zero energy [Eq. (3)], implying that neither the quantities nor the state evolve at all.
@@
In contrast to the MS setup, the quantum correlations and state dynamics for the NFS case exhibit an oscillatory behaviour with time, which start from unity reflecting the initial maximally entangled QDs, see Fig. 2(b). The distinct time evolution occurs because the contributing eigenvalues in Eq. (7) are finite for the NFS, implying that all the coefficients from Eqs. (25) contribute to the time evolution that becomes oscillatory over time. As time progresses, we observe that there are times where $C=0$ but $D\neq 0$, signalling that there is a finite quantum correlation between the two QDs even when the NFS reaches its separable state whose specific form is not evident. Also, Fig. 2(b) shows that there are other times $t$ at which $C=1$ and $D=1$ when $R_{p}=0$, a regime showing that additional maximally entangled states of the QDs are being generated during the dynamics in the NFS  [^1] At $t=(\pi/\sqrt{2})\,\hbar/\lambda_{1}$, the coefficients $\bar{c}_{2}(t)$ and $\bar{c}_{3}(t)$ are zero and $\bar{c}_{4}(t)=-1$; therefore, the state of the system becomes (|$(\ket{000}-\ket{110})/2$, which is also a maximally entangled state of the QDs, thus giving $C=1$.. For the return probability and entanglement dynamics of the NFS, their oscillatory behaviour can be easily seen analytically from Eqs. (26): $R_{p}=|(1+\cos{\sqrt{2}t})/2|^{2}$ with period $T=((2n+1)\pi/\sqrt{2})\,\hbar/\lambda_{1}$, being $n$ an integer and $\hbar/\lambda_{1}$ the unit of time, while $E_{d}=|(1+i\sqrt{2}\sin{\sqrt{2}t})/2|^{2}$ with period $T=(\pi/\sqrt{2})\,\hbar/\lambda_{1}$. It is straightforward to get $R_{p}=1$ and $E_{d}=1/4$ at $t=0$, while at $t=({n\pi}/{\sqrt{2}})\,\hbar/\lambda_{1}$ we get $R_{p}=0$ and $E_{d}$ develops minima leading to a fall of the state into the eigenstate |$\ket{000}$, see by green and black curves in Fig. 2(b). The overall behaviour of the entanglement measures in the NFS setup is clearly different to what we found for the MS system. Therefore, zero-energy MBSs and zero-energy regular normal fermions induce distinct entanglement signatures
@@
In the case of finite energies $\omega\neq 0$ at symmetric couplings $\lambda=1$, the time evolution of the entanglement measures exhibit a rather similar behaviour as the one for zero energy discussed in the previous two paragraphs, see Fig. 2(c,d). For the MS in Fig. 2(c), we find $R_{d}$, $C$, and $D$ exhibiting a constant value equal to $1$, while $E_{d}=0.25$, which is the same to what we saw in Fig. 2(a) at $\omega=0$. To understand the intriguing behaviour at $\omega\neq 0$, we note that the coefficient $\chi$ [Eq. (19)] of the time evolved state given by Eq. (18) vanishes at $\lambda=1$ while the other coefficient becomes $\eta(t)={\rm cos}(|\omega|t)+i{\rm sin}(|\omega|t)$ and is thus entirely determined by the energy splitting of MBSs. As a result, the time-evolved state from Eq. (18) evolves with a constant phase, determined by $\omega$, where $|\eta(t)|^{2}=1$ enables $C$, $D$, and $R_{p}$ to remain constant at $1$ and $E_{d}=0.25$, see Eq. (19). The quantities equal to $1$ then imply that the system remains with maximally entangled QDs and $E_{d}=0.25$ that there is only a contribution from the |$\ket{000}$ configuration in the entanglement dynamics. For the NFS at $\lambda=1$ and $\omega\neq 0$ in Fig. 2(d), the entanglement measures exhibit a time evolution that is similar to the $\omega=0$ case but with some subtle differences. At $\omega\neq 0$, the periodic $C$ and $D$ develop maxima, which, however, are not accompanied by $R_{p}=0$, implying that there is no revival of the initial state at short times  [^2] We have verified that the revival of the initial state occurs over a longer period of time than shown in the plot.. This situation implies that another entangled state, having a component of the initial state, is being generated during the evolution. Moreover, the minima of $E_{d}$ states that one of the configurations in |$\ket{\phi}$ remains in the state function throughout the dynamics.
@@
When the couplings become asymmetric $\lambda\neq 1$, the entanglement measures of the MS as well as those of the NFS develop an oscillatory profile, see Fig. 2(e-h). At $\omega=0$, the MS with zero-energy MBSs achieves $D=0$ and $C=0$ at certain times, which is contrary to the NFS case, where $D\neq 0$ when $C=0$, see Fig. 2(e,f). Besides this point, we also note that in the MS system the concurrence and discord reach unity when $R_{p}=0$, showing an another state, (|$(\ket{000}-\ket{110})/\sqrt{2}$, of two maximally entangled QDs has been created. Moreover, the entanglement dynamics for the MS remains constant because the initial state changes only to (|$(\ket{000}-\ket{110})/\sqrt{2}$, and then it revives back, resulting in a constant value of $E_{d}=0.25$, while it oscillates for the NFS, again contrasting the distinct signatures of both systems [Fig. 2(e,f)]. At finite energies ($\omega\neq 0$) and asymmetric couplings ($\lambda\neq 1$), the MS and the NFS exhibit similar properties as seen in Fig. 2(g,h) but still with some small differences. The similar behavior is associated to the oscillatory profile which stems from having extra configurations participating in the dynamics because the coefficient of the time-evolved state in Eq. (19) is non zero, $\chi\neq 0$. It is worth noting that the entanglement measures for the MS setup have larger periodicity [Fig. 2(g)], which occurs because the eigenenergies in Eq. (4) have contributions from $\lambda$ and $\omega$, see also Eqs. (19) and Eq. (18). Among the differences at $\omega\neq 0$ and $\lambda\neq 1$, we find that the Majorana system MS achieves regimes with $C$, $D$, and $R_{p}$ equal to $1$ (as shown in Fig. 2(g)), which means that the finite $\omega$ prefers the revival of the initial state rather than changing the local phase, which would result in achieving another maximally entangled state of quantum dots, as depicted for the MS in Fig. 2(e). This phenomenon, however, is not observed in the NFS, as shown in Fig. 2(h). Another difference is that in the NFS case, the quantities $E_{d}$ and $R_{p}$ exhibit a fall that reveal the presence of arbitrary configurations in the state; their peaks show that the different entangled states between the two QDs are being generated and destroyed. Thus, even though the oscillatory profile of the entanglement measures for the MS and NFS exhibit some similarities, there still are significant differences that could allow identifying MBSs.

### v2

@@ III Results for a maximally entangled initial state
In this part, we follow the discussion of previous section and obtain the quantum correlations and state dynamic probabilities for the two systems described in Section II, taking into account a maximally entangled initial state. We remind that the first introduced setup corresponds to a Majorana system (MS), where QDs are coupled via nonlocal fermion of Majorana origin. The second setup consists of QDs coupled via a normal fermion, coined normal fermion system (NFS). Since the manipulation of QDs was shown to be a feasible task [86, 87], see also Refs. [88, 89, 90], we consider that the two QDs are in a maximally entangled initial state, while the nonlocal fermion due to MBSs is in the zero state. We note that it is necessary to consider the zero state of the nonlocal fermion in order to maintain the even parity of the system, which is the sector we investigate; see Section II. The separability of the non-local fermion with entangled QDs is justified by assuming that the topological superconductor and QD are not coupled before the evolution of the entire system starts.
@@
where |$\ket{E_{j}}$ represents the eigenstates of the MS or NFS (|$(\ket{E_{j}}\rightarrow\ket{E_{j}^{\prime}})$ given by Eqs. (4) or Eqs. (7), depending on the system under investigation. Note that the initial state |$\ket{\psi(0)}$ is given in Eq. (17). Before going further, we point out that the contribution from the eigenstates in Eq. (18) also depends on the form of the initial states. In this regard, we have that the primary contributions in the MS come from eigenstates |$\ket{E_{1}}$ and |$\ket{E_{2}}$ with energies $-\Delta$ and $\Delta$, respectively. The eigenstates |$\ket{E_{3}}$ and |$\ket{E_{4}}$ are orthogonal to the initial state, $\left\langle E_{j}|\psi(0)\right\rangle=0$, which implies that they do not contribute to the dynamics given by Eq. (18). Contrary to the MS, the dynamics of the NFS is determined by the contributions of all states.
@@ III.1 Majorana system
Here, $\Delta$, $\lambda$, and $\omega$ represent the lowest positive eigenvalue, the coupling between nonlocal fermions and the right QD, and the energy splitting between MBSs, as described in Subsection 4. The coefficients $\eta$ and $\chi$ in the above state function shape the dynamics of the entanglement generations between the MBSs and the second QD. They can also be manipulated to see the type of maximally entangled state (of the Bell’s state form) that may be achieved in the evolution. Therefore, below, we now compute the quantities discussed in Section II.3.
@@ III.1.1 Return probability and entanglement dynamics
By using Eq. (16), we obtain the entanglement dynamics $E_{d}$ and the return probability $R_{p}$. For $E_{d}$ we need to specify the desired state onto which will be projected the time-evolved function |$\ket{\psi(t)}$. For this purpose, we consider |$\ket{\phi}$ describing a maximally entangled nonlocal fermion and the second QD given by
@@ III.2 Normal fermion system
In this part, we carry out the same calculations as in the previous subsection but when the QDs are coupled via a normal fermion. The time-evolved state function for the initial state from Eq. (17) is here given by
@@
$\begin{split}R_{p}&=\frac{|\bar{c}_{1}(t)+\bar{c}_{4}(t)|^{2}}{2}\,,\\ E_{d}&=\frac{|\bar{c}_{1}(t)+\bar{c}_{2}(t)|^{2}}{2}\,,\\ C&=2(||\bar{c}_{1}(t)||\bar{c}_{4}(t)|-|\bar{c}_{2}(t)||\bar{c}_{3}(t)||)\,,\end{split}$ (27)
@@
III.3 Discussion of state dynamics and quantum correlations
@@
Having presented the general expressions obtained of concurrence (C), discord (D), entanglement dynamics ($E_{d}$) and return probability ($R_{p}$) for the MS and the NFS with maximally entangled QDs, here we discuss their time evolution. To vizualize the time evolution, in Fig. 2 we plot C, D, $E_{d}$, and $R_{p}$ as a function of time $t$ for the two systems under consideration (MS and NFS). Fig. 2(a-d) and Fig. 2(e-h) correspond to QDs with symmetric ($\lambda=1$) and asymmetric ($\lambda\neq 1$) couplings in the MS or NFS. As discussed before, in the MS setup, QDs are coupled via MBSs, while QDs are coupled via a normal fermion in the NFS. To truly exploit the inherent Majorana nonlocality, in Fig. 2 we contrast regimes at zero and finite energies.
@@
We first analyze the case of symmetric couplings at zero energy, which corresponds to $\omega=0$ and $\lambda=1$ in Fig. 2(a,b). In the case of the MS, this regime has zero-energy MBSs and the first feature to notice is that the quantities $C$, $D$, and $R_{p}$ remain at $1$, while $E_{d}=0.25$, as shown in Fig. 2(a), see also subsection III.1. Since the initial state involves maximally entangled QDs, having a constant time evolution reveals that such a maximal entanglement is not affected. The constant time evolution can be understood by noting that the eigenstates of the Hamiltonian given by Eq. (4), except for the third eigenstate, are orthogonal to the initial state given by Eq. (18), giving a vanishing overlap that does not contribute to the dynamics as seen in Fig. 2(a). The third eigenstate of the MS is the same as the initial state [Eq. (17)] and has zero energy [Eq. (3)], implying that neither the quantities nor the state evolve at all.
@@
In contrast to the MS setup, the quantum correlations and state dynamics for the NFS case exhibit an oscillatory behaviour with time, which starts from unity, reflecting the initial maximally entangled QDs (see Fig. 2(b)). The distinct time evolution occurs because the contributing eigenvalues in Eq. (7) are finite for the NFS, implying that all the coefficients from Eqs. (26) contribute to the time evolution that becomes oscillatory over time. As time progresses, we observe that there are times where $C=0$ but $D\neq 0$, signalling that there is a finite quantum correlation between the two QDs even when the NFS reaches its separable state, whose specific form is not evident. Also, Fig. 2(b) shows that there are other times $t$ at which $C=1$ and $D=1$ when $R_{p}=0$, a regime showing that additional maximally entangled states of the QDs are being generated during the dynamics in the NFS  [^1] At $t=(\pi/\sqrt{2})\,\hbar/\lambda_{1}$, the coefficients $\bar{c}_{2}(t)$ and $\bar{c}_{3}(t)$ are zero and $\bar{c}_{4}(t)=-1$; therefore, the state of the system becomes (|$(\ket{000}-\ket{110})/2$, which is also a maximally entangled state of the QDs, thus giving $C=1$.. For the return probability and entanglement dynamics of the NFS, their oscillatory behaviour can be easily seen analytically from Eq. (27): $R_{p}=|(1+\cos{\sqrt{2}t})/2|^{2}$ with a period $T=\sqrt{2}\pi\,\hbar/\lambda_{1}$, where $\hbar/\lambda_{1}$ is the unit of time. Similarly, the entanglement dynamics is given by $E_{d}=|(1+i\sqrt{2}\sin{\sqrt{2}t})/2|^{2}$ with period $T=(\pi/\sqrt{2})\,\hbar/\lambda_{1}$. It is straightforward to get $R_{p}=1$ and $E_{d}=1/4$ at $t=0$, while at $t=({n\pi}/{\sqrt{2}})\,\hbar/\lambda_{1}$ we get $R_{p}=0$ and $E_{d}$ develop minima, leading to a fall of the state into the eigenstate |$\ket{000}$, as depicted by the green and black curves in Fig. 2(b). The overall behaviour of the entanglement measures in the NFS setup is clearly different to what we found for the MS. Therefore, zero-energy MBSs and zero-energy normal fermion induce distinct entanglement signatures.
@@
In the case of finite energies $\omega\neq 0$ at symmetric couplings $\lambda=1$, the time evolution of the entanglement measures exhibits a rather similar behaviour as the one for zero energy discussed in the previous two paragraphs; see Fig. 2(c,d). For the MS in Fig. 2(c), we find $R_{p}$, $C$, and $D$ exhibiting a constant value equal to $1$, while $E_{d}=0.25$, which is the same as what we saw in Fig. 2(a) at $\omega=0$. To understand the intriguing behaviour at $\omega\neq 0$, we note that the coefficient $\chi$ [Eq. (20)] of the time evolved state given by Eq. (19) vanishes at $\lambda=1$ while the other coefficient becomes $\eta(t)={\rm cos}(|\omega|t)+i{\rm sin}(|\omega|t)$ and is thus entirely determined by the energy splitting of MBSs. As a result, the time-evolved state from Eq. (19) evolves with a constant phase, determined by $\omega$, where $|\eta(t)|^{2}=1$ enables $C$, $D$, and $R_{p}$ to remain constant at $1$ and $E_{d}=0.25$, see Eq. (20). The quantities equal to $1$ then imply that the system remains with maximally entangled QDs and $E_{d}=0.25$ that there is only a contribution from the configuration |$\ket{000}$ in the entanglement dynamics. For the NFS at $\lambda=1$ and $\omega\neq 0$ in Fig. 2(d), the entanglement measures exhibit a time evolution that is similar to the $\omega=0$ case but with some subtle differences. At $\omega\neq 0$, the periodic $C$ and $D$ develop maxima, which, however, are not accompanied by $R_{p}=0$, implying that there is no revival of the initial state at short times  [^2] We have verified that the revival of the initial state occurs over a longer period of time than shown in the plot.. This situation implies that another entangled state, having a component of the initial state, is being generated during the evolution. Moreover, the minima of $E_{d}$ states that one of the configurations in |$\ket{\phi}$ remains in the state function throughout the dynamics.
When the couplings become asymmetric, $\lambda\neq 1$, the entanglement measures of the MS as well as those of the NFS develop an oscillatory profile; see Fig. 2(e-h). At $\omega=0$, the MS with zero-energy MBSs achieves $D=0$ and $C=0$ at certain times, which is contrary to the NFS case, where $D\neq 0$ when $C=0$, see Fig. 2(e,f). The evolution of the MS in Fig. 2(e) starts from the state $(|000\rangle+|110\rangle)/\sqrt{2}$, which has maximally entangled QDs maintaining the even parity of the entire system. Since the system is initially in a maximally entangled state between the two QDs, it transitions from maximal entanglement by reducing the contribution of the configuration |$\ket{110}$ and moves towards the separable state |$\ket{000}$. The system’s state is well-characterized as the entanglement dynamics appropriately yield $E_{d}=1/4$, due to the contribution solely from the configuration |$\ket{000}$. As the evolution progresses beyond this point, the configuration |$\ket{110}$ acquires a finite phase that eventually becomes $-1$; this process then results in the state $(|000\rangle-|110\rangle)/\sqrt{2}$, which is again a state with maximally entangled QDs characterized by $C=1$, $D=1$, $R_{p}=0$, and $E_{d}=1/4$.
It is important to note that the dynamics beyond the point when the system reaches |$\ket{000}$ involves only the additional configuration |$\ket{110}$, as other possibilities such as |$\ket{011}$ and |$\ket{101}$, are excluded. In fact, the configuration |$\ket{011}$ does not contribute because $E_{d}$ remains constant, and |$\ket{101}$ does not contribute because, physically, this would imply the generation of entanglement between the first QD and the nonlocal fermion from the MBSs, which would decrease the entanglement between the two QDs due to the monogamy property of entanglement. However, since the dynamics indicates increasing entanglement beyond the time when the system reaches |$\ket{000}$, we can be sure that only |$\ket{110}$ is present in the dynamical state of the system.
@@
Contrasting the constant value of $E_{d}=0.25$, it oscillates for the NFS, giving distinct signatures of both systems [Fig. 2(e,f)]. At finite energies ($\omega\neq 0$) and asymmetric couplings ($\lambda\neq 1$), the MS and the NFS exhibit similar properties as seen in Fig. 2(g,h) but still with some small differences. Similar behavior is associated with the oscillatory profile, which stems from having extra configurations participate in the dynamics because the coefficient of the time-evolved state in Eq. (20) is non zero, $\chi\neq 0$. It is worth noting that the entanglement measures for the MS setup have larger periodicity [Fig. 2(g)], which occurs because the eigenenergies in Eq. (4) have contributions from $\lambda$ and $\omega$; see also Eqs. (20) and Eq. (19). Among the differences at $\omega\neq 0$ and $\lambda\neq 1$, we find that the Majorana system MS achieves regimes with $C$, $D$, and $R_{p}$ equal to $1$ (as shown in Fig. 2(g)), which means that the finite $\omega$ prefers the revival of the initial state rather than changing the local phase, which would result in achieving another maximally entangled state of quantum dots, as depicted for the MS in Fig. 2(e). This phenomenon, however, is not observed in the NFS, as shown in Fig. 2(h). Another difference is that in the NFS case, the quantities $E_{d}$ and $R_{p}$ exhibit a fall that reveals the presence of arbitrary configurations in the state; their peaks show that the different entangled states between the two QDs are being generated and destroyed. Thus, even though the oscillatory profiles of the entanglement measures for the MS and NFS exhibit some similarities, there are still significant differences that could allow identifying the MBSs.
The characteristics of MS in $\omega\neq 0$ provide further insight into the distinguishing signatures of MBSs and ABSs. It is important to emphasize that MBSs form a nonlocal fermion that is sensitive to changes in the length of the superconductor. In contrast, ABSs often have finite energies and do not depend on the length of the system; ABSs are not nonlocal. As a result, the finite energy splitting $\omega$ mimics the behavior of ABS energy. As shown in Fig. 2, the correlation dynamics at $\omega\neq 0$ and asymmetric couplings exhibit an oscillatory time evolution that is similar for both MS and NFS, see Fig. 2(g,h). This supports the idea that ABSs, due to $\omega\neq 0$, are local states, much like a normal fermion. Thus, the consideration of $\omega=0$ and $\omega\neq 0$ highlights the distinct features of true zero-energy MBSs and ABSs, respectively.